Parth (parthkohli):

\[{d \over dx} e^{12w} = 12e^{12w - 1}w\]Did I do this right?

OpenStudy (lgbasallote):

no.. if you're deriving it with respect to x..then that's just a constant

Parth (parthkohli):

So 0?

OpenStudy (lgbasallote):

yup

Parth (parthkohli):

Oh, so\[{d \over de} e^{12w} = \text{ whatever I wrote there above}\]

OpenStudy (lgbasallote):

uhh still no... e is not your dependent variable

OpenStudy (lgbasallote):

should be \[\frac{d}{dw}\]

OpenStudy (lgbasallote):

and that's still not the derivative... \[\large \frac{d}{dw} e^{12w} \implies e^{12w} \times \frac{d}{dx} (12 w)\]

OpenStudy (lgbasallote):

uhh that should be d/dw on the last part there

Parth (parthkohli):

Oh, well...:|

Parth (parthkohli):

Thank you, lgba, my faith in you has been restored :)

OpenStudy (lgbasallote):

lol what o.O

Parth (parthkohli):

Never mind =)